A father is 32 years older than his son. In 4 years the father will be 5 times as old as his son. How old is each now?
Answer by sudhanshu_kmr(1152)


age of son = x
age of father = 32 +x
5(x+ 4)= (32 + x + 4)
=> 5x + 20 = 36 + x
=> 4x = 16
=> x = 4
age of son = 4
father's age = 36